the electron affinity of chlorine is 3.9ev how much energy in KCAl is released when 2g of chlorine is completely converted to cl- ion in gaseous state

Grade:12

1 Answers

Vikas TU
7 years ago
write the eqn.
Cl2        +       2e-        ----------->>>           2Cl-
2/71 moles                    would give             2/71*2 = > 1/71 moles.    
For 2 elctrons,
Energy = 3.9 * 2 /71 = >0.109 eV.
 = 0.109 * 1.6022 * 10^-22
 = 0.17676 * 10^-22 Kcalories.